### v1

# A Family of Convex Models to Achieve Fairness through Dispersion Control

Abstract
Controlling the dispersion of a subset of decision variables in an optimization problem is crucial for enforcing fairness or load-balancing across a wide range of applications. Building on the well-known equivalence of finite-dimensional norms, the note develops a family of parameterized convex models that regulate the dispersion of a vector of decision-variable values through its coefficient of variation. Each model contains a single parameter that takes a value in the interval $[0,1]$. When the parameter is set to zero, the model imposes only a trivial constraint on the optimization problem; when set to one, it enforces equality of all the decision variables. As the parameter varies, the coefficient of variation is provably bounded above by a monotonic function of that parameter. The note also presents theoretical results that relate the space of feasible solutions to all the models.
1 Introduction
Fairness has deep roots in game theory, where foundational bargaining solutions were rigorously characterized via axiomatic frameworks. Nash’s seminal work formalized the bargaining problem through a set of axioms–Pareto efficiency, symmetry, invariance to affine transformations, and independence of irrelevant alternatives–which uniquely identify the Nash bargaining solution nash1950bargaining . Subsequent refinements, such as the Kalai-Smorodinsky solution, altered these axioms to prioritize proportionality among agents kalai1977proportional . In economics and communication theory xinying2023guide ; shakkottai2008network , these game-theoretic fairness axioms guide the design of utility functions so that decentralized, self-interested agents naturally converge to equitable outcomes. By carefully embedding fairness criteria into each agent’s payoff, one ensures that the collective solution respects core notions of equity and parity, even when each participant acts independently.
However, in many application domains, problems are centrally optimized: a single decision maker minimizes a global cost or maximizes overall efficiency while imposing load-balancing requirements to distribute burdens–battery discharge, travel distance, computational load ghomi2017load ; janiar2021deep ; bhadoriya2024equitable –uniformly across agents. Here, “fairness” is more accurately viewed as dispersion control, wherein the vector of decision-variable values must exhibit limited spread under a centralized objective. In contemporary optimization, efforts have been made to extend beyond achieving minimal cost or maximal utility by explicitly regulating the dispersion of selected decision variables. By bounding or penalizing the spread through metrics such as variance, range, coefficient of variation, or more sophisticated equity indices xinying2023guide , one can embed fairness or load-balancing constraints directly into existing optimization models. However, many dispersion metrics do not admit convex reformulations; directly bounding them thus renders the optimization problem non‑convex and computationally challenging xinying2023guide . Nevertheless, the dispersion‑control paradigm is ubiquitous: ensuring equitable service levels in cloud-computing clusters ghomi2017load , allocating shared spectrum in wireless networks matyjas2016spectrum , operating power grids during extreme events sundar2025parametric , among others. Controlling dispersion thus provides a unifying formalism that couples classical efficiency criteria with societally critical notions of parity, equity, and fairness.
1.1 Literature Review
In the literature, two principal modelling paradigms have emerged for incorporating fairness into decision-making problems that traditionally focus on maximizing efficiency.
[…]
Fairness levels are typically assessed a posteriori using indices such as the Gini coefficient gini1936measure , Atkinson index atkinson1970measurement , Hoover index long1997hoover , and Jain’s index jain1984quantitative etc. A substantial body of work has sought to enforce fairness by directly minimizing or maximizing such indices, or by reducing dispersion measures such as the coefficient of variation or the range of decision variables xinying2023guide . These approaches, however, generally suffer from one of two drawbacks: (i) the resulting model is non-convex, or (ii) the model produces a single fair solution that compromises efficiency–an outcome often unacceptable in applications where cost considerations are critical. Despite these issues, this paradigm remains widely used, given the limited availability of alternative frameworks for fairness in centralized optimization.
The second, more recent paradigm–which motivates the present study–retains an efficiency-centric objective while constraining dispersion. Examples include the $\varepsilon$-fairness second-order cone approach sundar2025parametric or the imposition of explicit upper bounds on fairness indices such as the Gini coefficient bektacs2020using , Jain’s index, coefficient of variation or range xinying2023guide . Here, fairness is modelled by enforcing a lower (or upper) bound on the minimum (or maximum) acceptable fairness level while preserving the efficiency objective. This approach does not always yield a convex optimization problem, as noted in xinying2023guide , and even when convexity is preserved, computing a suitable lower bound for fairness often requires trial-and-error. Two notable exceptions are the $\varepsilon$-fairness second-order cone method sundar2025parametric , which achieves this using a single second-order cone constraint, and the direct bounding of the Gini coefficient bhadoriya2024equitable , which reformulates the problem into a set of linear inequalities at the cost of introducing $O(n^{2})$ auxiliary variables. For a comprehensive review of both paradigms, the reader is referred to the survey in xinying2023guide . This note examines $\varepsilon$-fairness and extends it to a family of convex models. The appeal of $\varepsilon$-fairness lies in its desirable properties: it enforces fairness through a single parametric convex constraint with $\varepsilon\in[0,1]$ as the parameter, and the parameter $\varepsilon$ is monotonic in the level of fairness. Moreover, it naturally facilitates the study of the Pareto front of optimal solutions balancing efficiency and fairness in a bi-objective setting bhadoriya2024equitable .
We now present the main contributions of this note in the context of this literature.
1.2 Contributions
The contributions of the note are as follows: we extend the recently introduced model of $\varepsilon$-fairness sundar2025parametric –that controls the dispersion of decision variables via its coefficient of variation everitt2010cambridge –to a countably infinite family of convex models. Each model is derived from the first principles through the classical equivalence of finite-dimensional norms halmos2017finite , namely between the $\ell_{p}$-norm and the $\ell_{1}$-norm for any $p\geqslant 2$. All models are governed by a single parameter, $\varepsilon\in[0,1]$: when $\varepsilon=0$, no constraint on the dispersion is imposed, and when $\varepsilon=1$, the constraint forces the dispersion to be zero. In particular, for $p=2$, the constraint reduces to a second-order cone sundar2025parametric ; for $p=\infty$, it reduces to a set of linear inequalities; and for $p\in(2,\infty)$, it reduces to the $\ell_{p}$-norm cone. In addition, we present theoretical results that characterize the relationships among these convex models for all values of $\varepsilon$, including their relative strength, inclusion properties, and points of equivalence.
2 Model Development
We begin by introducing notation and recalling a classical result on the equivalence of norms in finite-dimensional spaces halmos2017finite . Let $\bm{x}\in\mathbb{R}^{n}_{\geqslant 0}$ denote an $n$-dimensional non-negative vector, whose dispersion we wish to regulate. For any integer $p\geqslant 1$, define its $\ell^{p}$‐norm by
$\|\bm{x}\|_{p}\;=\;\Bigl(\sum_{i=1}^{n}x_{i}^{p}\Bigr)^{1/p},$ (1)
Here we omit the absolute value in (1) since $x_{i}\geqslant 0$ for all $i$. Also define
$D_{p}\;\triangleq\;n^{1-1/p}-1.$ (2)
Then, for any pair of integers $p_{1}$ and $p_{2}$, such that $0<p_{1}<p_{2}$, the norm‐equivalence result halmos2017finite states
$\mathcal{NE}(p_{1},p_{2}):\;\;\|\bm{x}\|_{p_{2}}\;\leqslant\;\|\bm{x}\|_{p_{1}}\;\leqslant\;\left(\frac{D_{p_{2}}+1}{D_{p_{1}}+1}\right)\,\|\bm{x}\|_{p_{2}}.$ (3)
Specializing to the case $(p_{1},p_{2})=(1,p)$ with $p\geqslant 2$, we obtain
$\mathcal{NE}(1,p):\;\;\|\bm{x}\|_{p}\;\leqslant\;\|\bm{x}\|_{1}\;\leqslant\;(D_{p}+1)\,\|\bm{x}\|_{p}.$ (4)
Moreover:
  • The lower bound $\|\bm{x}\|_{p}=\|\bm{x}\|_{1}$ is attained if and only if all but one component of $\bm{x}$ are zero (the most unfair distribution of $\bm{x}$).
  • The upper bound $\|\bm{x}\|_{1}=(D_{p}+1)\,\|\bm{x}\|_{p}$ is attained if and only if all components of $\bm{x}$ are equal (the most fair distribution of $\bm{x}$).
This intuition naturally leads to the following definition, extending the concept of $\varepsilon$-fairness in sundar2025parametric .
Definition 1
$(\varepsilon,p)$‐fairness – For any integer $p\geqslant 2$, $\bm{x}\in\mathbb{R}^{n}_{\geqslant 0}$ and $\varepsilon\in[0,1]$, we say that $\bm{x}$ is $(\varepsilon,p)$‐fair if
$\bigl(1+\varepsilon\,D_{p}\bigr)\,\|\bm{x}\|_{p}\;=\;\|\bm{x}\|_{1}.$
Furthermore, we say that $\bm{x}$ is at least $(\varepsilon,p)$‐fair if
$\bigl(1+\varepsilon\,D_{p}\bigr)\,\|\bm{x}\|_{p}\;\leqslant\;\|\bm{x}\|_{1}.$ (5)
Additionally, we denote the set of at least $(\varepsilon,p)$-fair vectors as
$\mathcal{X}(\varepsilon,p)\;\triangleq\;\bigl\{\bm{x}\in\mathbb{R}_{\geqslant 0}^{n}:(1+\varepsilon\,D_{p})\,\|\bm{x}\|_{p}\leqslant\|\bm{x}\|_{1}\bigr\}.$ (6)
In the above definition, the factor $(1+\varepsilon D_{p})$ can be interpreted as the convex combination of $1$ and $(D_{p}+1)$, with $\varepsilon$ as the non-negative multiplier. Furthermore, the constraint in (5) is convex for every $p\geqslant 2$. In particular, for $p=2$, (5) reduces to a second‐order cone sundar2025parametric ; for $2<p<\infty$, it defines an $\ell_{p}$‐norm cone, and for $p=\infty$, it reduces to a collection of linear inequalities as follows:
$\|\bm{x}\|_{1}\;\geqslant\;(1+\varepsilon D_{\infty})\,\|\bm{x}\|_{\infty}\iff\sum_{i=1}^{n}x_{i}\;\geqslant\;\left(1+n\varepsilon-\varepsilon\right)\cdot x_{i}\quad\forall\,i.$ (7)
For any fixed $\varepsilon$, the Definition 1 yields a countably infinite family of convex models (by varying the integer $p\geqslant 2$) for enforcing fairness in the distribution of $\bm{x}$. We now present a proposition that directly follows from the definition of $\mathcal{X}(\varepsilon,p)$ in (6).
Proposition 1
For any integer $p\geqslant 2$ and and $\varepsilon_{1},\varepsilon_{2}\in[0,1]$ such that $\varepsilon_{1}>\varepsilon_{2}$, $\mathcal{X}(\varepsilon_{1},p)\subset\mathcal{X}(\varepsilon_{2},p)$. ∎
The following normalization argument to simplifies notation for the main theoretical results in the subsequent section.
Remark 1
Normalization – Since all finite-dimensional norms are positively homogeneous of degree 1, the inequality in (6), i.e.,
$\bigl(1+\varepsilon\,D_{p}\bigr)\,\|\bm{x}\|_{p}\;\leqslant\;\|\bm{x}\|_{1}$
is invariant under scaling of $\bm{x}$. Thus, it is convenient to normalize any non-zero vector [^1] The zero vector, $\bm{0}$, cannot be normalized and is trivially $(\varepsilon,p)$-fair for any $p\geqslant 2$ and $\varepsilon\in[0,1]$. Hence, throughout the rest of the note, we assume that $\bm{x}$ is always a non-zero vector. by its $\ell_{1}$-norm. To that end, define
$\Delta_{n}\;\triangleq\;\left\{\bm{x}\in\mathbb{R}_{\geqslant 0}^{n}:\sum_{i=1}^{n}x_{i}=1\right\},$ (8)
the standard probability simplex in $\mathbb{R}^{n}$. Normalizing any $\bm{x}\in\mathcal{X}(\varepsilon,p)$ with respect to its $\ell_{1}$-norm then yields
$\mathcal{Y}(\varepsilon,p)\;\triangleq\;\left\{\bm{x}\in\Delta_{n}:(1+\varepsilon\,D_{p})\,\|\bm{x}\|_{p}\leqslant 1\right\}.$ (9)
Observe that $\mathcal{Y}(\varepsilon,p)$ is exactly the image of $\mathcal{X}(\varepsilon,p)$ under the map $\bm{x}\mapsto\bm{x}/(\bm{e}^{\intercal}\bm{x})$; here $\bm{e}$ is a vector of 1s. Hence, any theoretical result for $\mathcal{X}(\varepsilon,p)$ can be transferred immediately to $\mathcal{Y}(\varepsilon,p)$ and vice-versa. All theoretical properties of $(\varepsilon,p)$-fairness in the subsequent section will be proved for the normalized set $\mathcal{Y}(\varepsilon,p)$.
3 Theoretical Properties
We first establish the relationship between the $(\varepsilon,p)$-fairness of Definition 1 and the coefficient of variation of the vector $\bm{x}$. The coefficient of variation of $\bm{x}$ is a statistical measure of the dispersion of $\bm{x}$. The result in the subsequent section 3.1 provides the intuition for how varying $\varepsilon$ in (5) for any fixed integer $p\geqslant 2$ allows the control of dispersion of $\bm{x}$.
3.1 Bound of Coefficient of Variation
The first property we establish is that the $(\varepsilon,p)$‐fairness constraint (5) implies an upper bound on the coefficient of variation of $\bm{x}$, and that this bound tightens as $\varepsilon$ increases.
Proposition 2
For any $\varepsilon\in[0,1]$, an integer $p\geqslant 2$, and $\bm{x}\in\mathcal{Y}(\varepsilon,p)$ defined in (9), the coefficient of variation of $\bm{x}$
$\mathrm{CV}(\bm{x})\triangleq\frac{\sqrt{\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\mu)^{2}}}{\mu},\quad\text{ where }\quad\mu=\frac{1}{n},$
satisfies the following bound:
$\left(\mathrm{CV}(\bm{x})\right)^{2}\;\leqslant\;B_{p}(\varepsilon)\;\triangleq\;\frac{(D_{p}+1)^{2}}{(1+\varepsilon\,D_{p})^{2}}-1\,.$ (10)
Moreover, $B_{p}(\varepsilon)$ is strictly decreasing in $\varepsilon$ on $[0,1]$, with $B_{p}(0)=(D_{p}+1)^{2}-1$ and $B_{p}(1)=0$.
Proof
For any $p\geqslant 2$, using norm equivalence $\mathcal{NE}(2,p)$ in (3) and (9), we have
$\|\bm{x}\|_{2}\;\leqslant\;\frac{D_{p}+1}{\sqrt{n}}\|\bm{x}\|_{p}\;\leqslant\;\frac{D_{p}+1}{\sqrt{n}}\left(\frac{1}{1+\varepsilon\,D_{p}}\right).$ (11)
On the other hand, since $n\mu=1$ and observing that
$\frac{1}{n}\sum_{i=1}^{n}x_{i}^{2}=\frac{\|\bm{x}\|_{2}^{2}}{n},\quad\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\mu)^{2}=\frac{\|\bm{x}\|_{2}^{2}}{n}-\mu^{2},$
we obtain
$\left(\mathrm{CV}(\bm{x})\right)^{2}=\frac{\|\bm{x}\|_{2}^{2}-n\mu^{2}}{n\mu^{2}}=n\,\|\bm{x}\|_{2}^{2}-1.$
Substituting the bound on $\|\bm{x}\|_{2}$ from (11) yields
$n\,\|\bm{x}\|_{2}^{2}-1$ $\;\leqslant\;\frac{(D_{p}+1)^{2}}{(1+\varepsilon\,D_{p})^{2}}-1.$
Hence,
$\left(\mathrm{CV}(\bm{x})\right)^{2}\;\leqslant\;\frac{(D_{p}+1)^{2}}{(1+\varepsilon\,D_{p})^{2}}-1=B_{p}(\varepsilon).$
Finally, since $D_{p}>0$, we have
$B_{p}^{\prime}(\varepsilon)\;=\;-\,\frac{2D_{p}(D_{p}+1)^{2}}{(1+\varepsilon\,D_{p})^{3}}\;<\;0.$
Hence, $B_{p}(\varepsilon)$ is strictly decreasing in $\varepsilon$ and $B_{p}(0)=(D_{p}+1)^{2}-1$, $B_{p}(1)=0$, as claimed.
The strict monotonicity of $B_{p}(\varepsilon)$ in $\varepsilon$ implies that, as $\varepsilon$ increases, the upper bound on $\mathrm{CV}(\bm{x})$ decreases, thereby decreasing the dispersion of $\bm{x}$. This relationship was previously established for the case $p=2$ in sundar2025parametric .
In the forthcoming sections, we present the main theoretical results that concern the equivalence or ordered relation between sets $\mathcal{Y}(p,\varepsilon)$ and correspondingly of $\mathcal{X}(p,\varepsilon)$ for varying values of $p$.
3.2 Equivalence Condition
In this section, we give conditions under which requiring $\bm{x}$ to be at least $(\varepsilon,p)$‑fair is equivalent for every integer $p\geqslant 2$. We start with a key result that directly follows from the definition of $(\varepsilon,p)$-fairness.
Proposition 3
Suppose that $\bm{e}_{i}$ is a unit vector with the $i$th coordinate set to $1$ and $\bm{e}$ is a vector of 1s, then for any integer $p\geqslant 2$, the following statements hold:
$\text{For any }i\in\{1,\dots,n\},~~\bm{e}_{i}\in\mathcal{Y}(\varepsilon,p)~\Longrightarrow~\varepsilon=0$ (12)
 $\text{ and }\varepsilon=1\iff\mathcal{Y}(\varepsilon,p)=\left\{\tfrac{1}{n}\bm{e}\right\}.$ (13)
Proof
The proof follows from the observation in Section 2 that the term $(1+\varepsilon D_{p})$ is the convex combination of the two scalars $1$ and $(D_{p}+1)$ with $\varepsilon$ as the factor.
Theorem 3.1
For any integer $p\geqslant 2$, and $\varepsilon\in\{0,1\}$, the sets of at least $(\varepsilon,p)$-fair vectors are equivalent, i.e.
$\mathcal{Y}(\varepsilon,p)=\mathcal{Y}(\varepsilon,2)\quad\forall~p>2.$
Proof
We treat the two cases $\varepsilon=0$ and $\varepsilon=1$ separately. When $\varepsilon=0$, the inequality in (9) is trivially satisfied for any $\bm{x}\in\Delta_{n}$ and any $p\geqslant 2$. Hence,
$\mathcal{Y}(0,p)=\Delta_{n}=\mathcal{Y}(0,2)\quad\forall,p>2.$
When $\varepsilon=1$, Proposition 3 indicates that $\mathcal{Y}(1,p)=\left\{\tfrac{1}{n}\bm{e}\right\}$ for every $p\geqslant 2$.
Theorem 3.1 shows that at the two extreme fairness levels, the feasible set of vectors $\bm{x}$ is the same regardless of the choice of $p$.
3.3 Strict Inclusion Property
Here, we prove a result that shows a strict inclusion property between $\mathcal{Y}(\varepsilon,p_{1})$ and $\mathcal{Y}(\varepsilon,p_{2})$ for any pair of integers $(p_{1},p_{2})$ such that $2\leqslant p_{1}<p_{2}$, and $\varepsilon$ in the open interval $(0,1)$. We begin by presenting a proposition that will be useful to prove the main theorem.
Proposition 4
For any $\bm{x}\in\Delta_{n}\setminus\left\{\bm{e}_{1},\dots,\bm{e}_{n},\tfrac{1}{n}\bm{e}\right\}$,
$F(p)\triangleq\frac{1-\|\bm{x}\|_{p}}{D_{p}\|\bm{x}\|_{p}}\text{ is strictly decreasing in }p\geqslant 2.$ (14)
Proof
See Appendix A.
Theorem 3.2
For any pair of integers $(p_{1},p_{2})$ such that $2\leqslant p_{1}<p_{2}$ and $\varepsilon\in(0,1)$, $\mathcal{Y}(\varepsilon,p_{2})\subset\mathcal{Y}(\varepsilon,p_{1})$.
Proof
For each $\bm{x}\in\Delta_{n}$ and integer $p\geqslant 2$, define
$\varepsilon_{p}(\bm{x})\;\triangleq\;\frac{1-\|\bm{x}\|_{p}}{D_{p}\,\|\bm{x}\|_{p}}.$ (15)
Since
$(1+\varepsilon\,D_{p})\,\|\bm{x}\|_{p}\leqslant 1\iff\varepsilon\leqslant\varepsilon_{p}(\bm{x})$
we have
$\bm{x}\in\mathcal{Y}(\varepsilon,p)\iff 0<\varepsilon\leqslant\varepsilon_{p}(\bm{x}).$
Therefore $\mathcal{Y}(\varepsilon,p_{2})\subset\mathcal{Y}(\varepsilon,p_{1})$ will follow once we show
$\varepsilon_{p_{1}}(\bm{x})>\varepsilon_{p_{2}}(\bm{x})~~\forall~~\bm{x}\in\Delta_{n}\setminus\left\{\bm{e}_{1},\dots,\bm{e}_{n},\tfrac{1}{n}\bm{e}\right\}.$ (16)
The excluded cases correspond to the boundary situation $\varepsilon\in\{0,1\}$, already treated in Proposition 3 and Theorem 3.1. For any such $\bm{x}\in\Delta_{n}\setminus\left\{\bm{e}_{1},\dots,\bm{e}_{n},\tfrac{1}{n}\bm{e}\right\}$, the proof is complete if we show that the map
$p\mapsto\varepsilon_{p}(\bm{x})=\frac{1-\|\bm{x}\|_{p}}{D_{p}\|\bm{x}\|_{p}}$
is strictly decreasing on $2,\infty]$ which is exactly the statement of Proposition 4. Figure 1 illustrates the idea graphically.
Figure 1: Graphical illustration for $\varepsilon_{p_{1}}(\bm{x})>\varepsilon_{p_{2}}(\bm{x})$.
Theorem 3.2 shows that the size of the feasible sets of requiring $\bm{x}$ to be at least $(\varepsilon,p)$-fair decreases with $p$ when $\varepsilon\in(0,1)$ and Theorem 3.1 shows that the feasible sets are equal for any value of integer $p\geqslant 2$ when $\varepsilon\in\{0,1\}$.
4 Conclusion
The note presents a countably infinite family of convex models–parameterized by $\varepsilon\in[0,1]$–to control the dispersion of a subset of decision variables in an optimization problem. The parameter $\varepsilon$ bounds dispersion: when $\varepsilon=0$, dispersion is unconstrained; when $\varepsilon=1$, dispersion is forced to zero. This family is indexed by integers $p\geqslant 2$, and all the models in the family are derived from first principles using the fundamental result of equivalence of finite-dimensional norms. For $p=2$, the model is a second‑order cone; for $2<p<\infty$, it is an $\ell_{p}$‑norm cone; and for $p=\infty$, it reduces to a set of linear inequalities. The note also establishes several theoretical properties that align with intuitive notions of fair solutions. The $p=\infty$ variant is appealing computationally since the model is linear. Finally, these models can be integrated into optimization models in many application domains to control dispersion or enforce the fairness of a vector of decision variables with minimal modifications to existing solution algorithms.
Appendix A Proof of Proposition 4
We start with introducing additional notation and a few lemmas that will make the presentation of the proof of the main result cleaner. For any vector $\bm{x}\in\Delta_{n}\setminus\left\{\bm{e}_{1},\dots,\bm{e}_{n},\tfrac{1}{n}\bm{e}\right\}$, we let
$S_{p}=\sum_{i=1}^{n}x_{i}^{p}~\Rightarrow~~S_{p}^{\frac{1}{p}}=\|\bm{x}\|_{p}$ (17)
and
$w_{i}=\frac{x_{i}^{p}}{S_{p}}~~\Rightarrow~~\sum_{i=1}^{n}w_{i}=1\text{ and }\sum_{i=1}^{n}w_{i}^{\frac{1}{p}}=\frac{1}{\|\bm{x}\|_{p}}.$ (18)
Also, applying $\mathcal{NE}(1,p)$ to $\bm{x}$ yields
$\|\bm{x}\|_{p}<1<(D_{p}+1)~\|\bm{x}\|_{p}~~\Longrightarrow~~\|\bm{x}\|_{p}\in\left((D_{p}+1)^{-1},1\right)$ (19)
Observe that $\bm{x}=\bm{e}_{i}$ for any $i\in\{1,\dots,n\}$ if and only if $\|\bm{x}\|_{p}=1$ and $\bm{x}=\tfrac{1}{n}\bm{e}$ if and only if $\|\bm{x}\|_{p}=(D_{p}+1)^{-1}$. Since these values are excluded from the possible values $\bm{x}$ can take, the inequality in (19) is strict.
Lemma 1
For any $\bm{x}\in\Delta_{n}\setminus\left\{\bm{e}_{1},\dots,\bm{e}_{n},\tfrac{1}{n}\bm{e}\right\}$ and integer $p\geqslant 2$, the following inequality holds
$\frac{p}{p-1}\left[\frac{-\ln\|\bm{x}\|_{p}}{1-\|\bm{x}\|_{p}}\right]-\ln n\left[1+\frac{1}{D_{p}}\right]<0$ (20)
Proof
We start the proof by observing that the function
$f(\|\bm{x}\|_{p})=\frac{-\ln\|\bm{x}\|_{p}}{1-\|\bm{x}\|_{p}}~~\text{ strictly decreases for $\|\bm{x}\|_{p}\in(0,1)$.}$
The above result is easy to see by evaluating the derivative and checking that the sign of the derivative is strictly negative. Now, for any fixed $p\geqslant 2$, we define $t_{0}=(D_{p}+1)^{-1}$. Since $\|\bm{x}\|_{p}\mapsto f(\|\bm{x}\|_{p})$ is a strictly decreasing function, for every value of $\|\bm{x}\|_{p}>t_{0}$ we have $f(\|\bm{x}\|_{p})<f(t_{0})$, i.e.,
$-\frac{\ln\|\bm{x}\|_{p}}{1-\|\bm{x}\|_{p}}$ $<-\frac{\ln(D_{p}+1)^{-1}}{1-(D_{p}+1)^{-1}}$
$=\frac{p-1}{p}~\ln n\left[1+\frac{1}{D_{p}}\right]$
$\Rightarrow~~\frac{p}{p-1}\left[\frac{-\ln\|\bm{x}\|_{p}}{1-\|\bm{x}\|_{p}}\right]$ $<\ln n\left[1+\frac{1}{D_{p}}\right]$
$\Rightarrow~~\frac{p}{p-1}\left[\frac{-\ln\|\bm{x}\|_{p}}{1-\|\bm{x}\|_{p}}\right]$ $-\ln n\left[1+\frac{1}{D_{p}}\right]<0.$
We remark that when $\bm{x}=\tfrac{1}{n}\bm{e}$, $\|\bm{x}\|_{p}$ evaluates to $t_{0}$ and the above inequality is satisfied at equality, and the claim of the lemma does not hold. Hence, $\bm{x}\neq\tfrac{1}{n}\bm{e}$ is assumed in the statement of the lemma.
Lemma 2
Suppose $\bm{w}=[w_{1},\dots,w_{n}]^{\intercal}$, then for any $p\geqslant 2$, the function $H(\bm{w})=-\sum_{i}w_{i}\ln w_{i}$ satisfies
$0\leqslant H(\bm{w})\leqslant-\frac{p}{p-1}\ln\|\bm{x}\|_{p}.$ (21)
Proof
The definitions in (18) tell us that $\bm{w}=[w_{1},\dots,w_{n}]^{\intercal}$ is a probability vector. Hence, $H(\bm{w})$ is the Shannon-entropy of $\bm{w}$, which we know is non-negative gallager1968information . To derive the upper bound, we use the definition of Rényi entropy of order $\alpha$ renyi1961measures given below
$H_{\alpha}(\bm{w})=\frac{1}{1-\alpha}\ln\sum_{i}w_{i}^{\alpha}.$ (22)
The following two facts about $H_{\alpha}(\cdot)$ are well known: $H_{1}(\bm{w})=H(\bm{w})$ and $H_{\alpha}(\cdot)$ is a decreasing function of $\alpha$. We now let $\alpha=\frac{1}{p}$. Then since $\alpha<1$ (because $p\geqslant 2$), we have
$H(\bm{w})\leqslant H_{\frac{1}{p}}(\bm{w})=\frac{p}{p-1}\ln\sum_{i=1}^{n}w_{i}^{\frac{1}{p}}=-\frac{p}{p-1}\ln\|\bm{x}\|_{p}.$ (23)
The final equality follows from (18).
Lemma 3
For any integer $p\geqslant 2$ and $\bm{x}\in\Delta_{n}$, given the definitions in (17), (18) and Lemma 2, the following relationship holds
$\ln S_{p}-p\frac{S_{p}^{\prime}}{S_{p}}=H(\bm{w})$
where $S_{p}^{\prime}$ is the derivative of $S_{p}$ with respect to pp.
Proof
We start by computing the derivative of $S_{p}$ with respect to $p$ as
$S_{p}^{\prime}$ $=\sum_{i=1}^{n}x_{i}^{p}\ln x_{i}=S_{p}\sum_{i=1}^{n}w_{i}\ln x_{i}$
$\iff p\frac{S_{p}^{\prime}}{S_{p}}$ $=\sum_{i=1}^{n}w_{i}\ln x_{i}^{p}=\sum_{i=1}^{n}w_{i}\ln(S_{p}w_{i})$
$\iff p\frac{S_{p}^{\prime}}{S_{p}}$ $=\ln S_{p}\left(\sum_{i=1}^{n}w_{i}\right)+\sum_{i=1}^{n}w_{i}\ln w_{i}$
$=\ln S_{p}-H(\bm{w})$
$\iff H(\bm{w})$ $=\ln S_{p}-r\frac{S_{p}^{\prime}}{S_{p}}.$
We now present the proof of Proposition 4 by performing direct sign analysis of the derivative of $F(p)$. We start by taking logarithm of $F(p)$ as
$\ln F(p)\triangleq g(p)=\ln(1-\|\bm{x}\|_{p})-\ln\|\bm{x}\|_{p}-\ln D_{p}.$ (24)
We now let $F^{\prime}(p)$, $g^{\prime}(p)$, $\|\bm{x}\|_{p}^{\prime}$ and $D^{\prime}_{p}$ denote the derivative with respect to $p$ of $F(p)$, $g(p)$, $\|\bm{x}\|_{p}$ and $D_{p}$, respectively. Then, we can differentiate (24), we get
$\frac{F^{\prime}(p)}{F(p)}=g^{\prime}(p)=-\frac{\|\bm{x}\|_{p}^{\prime}}{1-\|\bm{x}\|_{p}}-\frac{\|\bm{x}\|_{p}^{\prime}}{\|\bm{x}\|_{p}}-\frac{D^{\prime}_{p}}{D_{p}}$
Simplifying the above equation further, we obtain
$g^{\prime}(p)$ $=-\frac{\|\bm{x}\|_{p}^{\prime}}{\|\bm{x}\|_{p}(1-\|\bm{x}\|_{p})}-\frac{D^{\prime}_{p}}{D_{p}}$
$=\frac{-1}{(1-\|\bm{x}\|_{p})}\left[\frac{1}{p}\frac{S_{p}^{\prime}}{S_{p}}-\frac{1}{p^{2}}\ln S_{p}\right]-\frac{\ln n}{p^{2}}\left[1+\frac{1}{D_{p}}\right]$
[…]
$\leqslant\frac{1}{p^{2}}\left(\frac{p}{p-1}\left[\frac{-\ln\|\bm{x}\|_{p}}{1-\|\bm{x}\|_{p}}\right]-\ln n\left[1+\frac{1}{D_{p}}\right]\right)$
[…]
The penultimate inequality follows from Lemma 2. We have proved that $g^{\prime}(p)<0$. But $F^{\prime}(p)=F(p)\cdot g^{\prime}(p)$ and since $F(p)>0$ for any $\bm{x}$ that satisfies the assumptions of the proposition, we obtain $F^{\prime}(p)<0$. ∎
Acknowledgements.
The authors thank Dirk Lauinger of the MIT Energy Initiative for insightful discussions on the potential for expanding on a family of fairness models.
Conflict of interest
The authors declare that they have no conflict of interest.