### v2

# A Family of Convex Models to Achieve Fairness through Dispersion Control

Abstract
Controlling the dispersion of a subset of decision variables in an optimization problem is crucial for enforcing fairness or load-balancing across a wide range of applications, from robotics to communication networks, supply-chain management, as well as public policy design. Building on the well-known equivalence of finite-dimensional norms, we develop a family of parameterized convex models that regulate the dispersion of a vector of decision-variable values through its coefficient of variation. Each model contains a single parameter defined on the interval $[0,1]$. When the parameter is set to 0, the model imposes only a trivial constraint on the optimization problem; when set to 1, it forces all decision variable values to become equal. As the parameter varies, the coefficient of variation is provably bounded above by a monotonic function of that parameter. We present theoretical results that relate the feasibility space of all the models and compare the models’ computational time and solution quality on a multi-agent assignment problem. Finally, we discuss the implications of these models for optimization problems in other domains.
I Introduction
Fairness has deep roots in game theory, where foundational bargaining solutions were rigorously characterized via axiomatic frameworks. Nash’s seminal work formalized the bargaining problem through a set of axioms–Pareto efficiency, symmetry, invariance to affine transformations, and independence of irrelevant alternatives–which uniquely identify the Nash bargaining solution [1]. Subsequent refinements, such as the Kalai-Smorodinsky solution, altered these axioms to prioritize proportionality among agents [2]. In economics and communication theory [3, 4], these game-theoretic fairness axioms guide the design of utility functions so that decentralized, self-interested agents naturally converge to equitable outcomes. By carefully embedding fairness criteria into each agent’s payoff, one ensures that the collective solution respects core notions of equity and parity, even when each participant acts independently.
However, in many application domains like control, robotics, and logistics, problems are centrally optimized: a single decision maker minimizes a global cost or maximizes overall efficiency while imposing load-balancing requirements to distribute burdens–battery discharge, travel distance, computational load [5, 6, 7]–uniformly across agents. Here, “fairness” is more accurately viewed as dispersion control, wherein the vector of decision-variable values must exhibit limited spread under a centralized objective.
In contemporary optimization, efforts have been made to extend beyond achieving minimal cost or maximal utility by explicitly regulating the dispersion of selected decision variables. By bounding or penalizing the spread through metrics such as variance, range, coefficient of variation, or more sophisticated equity indices [3], one can embed fairness or load-balancing constraints directly into existing optimization models. However, many dispersion metrics do not admit convex reformulations; directly bounding them thus renders the optimization problem non‑convex and computationally challenging [3]. Nevertheless, the dispersion‑control paradigm is ubiquitous: ensuring equitable service levels in cloud-computing clusters [5, 8], allocating shared spectrum in wireless networks [6, 9], managing electric-vehicle fleet loads [7], operating power grids during extreme events [10], and coordinating distributed energy resources [11], among others.
Controlling dispersion thus provides a unifying formalism that couples classical efficiency criteria with societally critical notions of parity, equity, and fairness. In the literature, two principal modeling paradigms have emerged. The first replaces the efficiency objective with an explicitly fairness-oriented objective–e.g., Rawlsian max–min fairness [12], $\alpha$-fairness [13], proportional fairness [4], or $\ell_{p}$-norm dispersion objectives [14]–so that parity is pursued as the primary goal. The second, which motivates the present study, retains an efficiency-centric objective while constraining dispersion, for example via the $\varepsilon$-fairness second-order cone [10] or explicit upper bounds on canonical indices such as the Gini coefficient [15, 14] and the Jain-index [16, 3]. The exhaustive survey in [3] reviews both approaches, and recent applications of the “efficiency-subject-to-fairness” paradigm span vehicle routing [7], power-grid operations and control [10, 11], and inventory control [17].
The contributions of the article are as follows: we extend the recently introduced model of $\varepsilon$-fairness [10]–that controls the dispersion of decision variables via its coefficient of variation [18]–to a countably infinite family of convex models. Each model is derived from the first principles through the classical equivalence of finite-dimensional norms [19], namely between the $p$-norm and the $1$-norm for any $p\geqslant 2$. All models are governed by a single parameter, $\varepsilon\in[0,1]$: when $\varepsilon=0$, no constraint on the dispersion is imposed, and when $\varepsilon=1$, the constraint forces the dispersion to be zero. In particular, for $p=2$, the constraint reduces to a second-order cone [10]; for $p=\infty$, it reduces to a set of linear inequalities; and for $p\in(2,\infty)$, it yields the familiar family of $\ell_{p}$-norm cones. The article also provides a comprehensive set of theoretical results that characterize the relationships among these convex models for all values of $\varepsilon$, detailing their relative strength, inclusion properties, and points of equivalence. Finally, we compare and contrast the different models in terms of computation time and its effectiveness in controlling dispersion on a multi-agent task assignment problem–an important problem that arises in many civilian and military applications like target tracking [20], intelligent transportation [21] and logistics [22], and search and rescue [23]. In particular, we formulate a variant of the assignment problem that aims to control the dispersion of the total time consumed to complete the tasks across different agents.
The rest of the article is organized as follows: Section II develops the fairness models from first principles, followed by theoretical properties of the models in Section III. Section IV formulates the fairness-constrained multi-agent assignment problem to demonstrate the use of the fairness models. Computational results that corroborate the effectiveness of the fairness models are presented in Section V followed by concluding remarks in Section VI.
II Model Development
We begin by introducing notation and recalling a classical result on the equivalence of norms in finite-dimensional spaces [19]. Let $\bm{x}\in\mathbb{R}^{n}_{\geqslant 0}$ denote an $n$-dimensional non-negative vector, whose dispersion we wish to regulate. For any integer $p\geqslant 1$, define its $\ell^{p}$‐norm by
$\|\bm{x}\|_{p}\;=\;\Bigl(\sum_{i=1}^{n}x_{i}^{p}\Bigr)^{1/p},$ (1)
Here we omit the absolute value in (1) since $x_{i}\geqslant 0$ for all $i$. Define also
$D_{p}\;\triangleq\;n^{1-1/p}-1.$ (2)
Then, for any pair of integers $p_{1}$ and $p_{2}$, such that $0<p_{1}<p_{2}$, the norm‐equivalence result [19] states
$\mathcal{NE}(p_{1},p_{2}):\;\;\|\bm{x}\|_{p_{2}}\;\leqslant\;\|\bm{x}\|_{p_{1}}\;\leqslant\;\frac{D_{p_{2}}+1}{D_{p_{1}}+1}\,\|\bm{x}\|_{p_{2}}.$ (3)
Specializing to the case $(p_{1},p_{2})=(1,p)$ with $p\geqslant 2$, we obtain
$\mathcal{NE}(1,p):\;\;\|\bm{x}\|_{p}\;\leqslant\;\|\bm{x}\|_{1}\;\leqslant\;(D_{p}+1)\,\|\bm{x}\|_{p}.$ (4)
Moreover:
  • The lower bound $\|\bm{x}\|_{p}=\|\bm{x}\|_{1}$ is attained if and only if all but one component of $\bm{x}$ are zero (the most unfair distribution).
  • The upper bound $\|\bm{x}\|_{1}=(D_{p}+1)\,\|\bm{x}\|_{p}$ is attained if and only if all components of $\bm{x}$ are equal (the most fair distribution).
This intuition naturally leads to the following definition, extending the concept of $\varepsilon$-fairness in [10].
Definition 1.
$(\varepsilon,p)$‐fairness – For any integer $p\geqslant 2$, $\bm{x}\in\mathbb{R}^{n}_{\geqslant 0}$ and $\varepsilon\in[0,1]$, we say that $\bm{x}$ is $(\varepsilon,p)$‐fair if
$\bigl(1+\varepsilon\,D_{p}\bigr)\,\|\bm{x}\|_{p}\;=\;\|\bm{x}\|_{1}.$
Thus, $\varepsilon=0$ corresponds to the most unfair distribution, and $\varepsilon=1$ to the perfectly fair distribution. Equivalently, we say that $\bm{x}$ is at least $(\varepsilon,p)$‐fair if
$\bigl(1+\varepsilon\,D_{p}\bigr)\,\|\bm{x}\|_{p}\;\leqslant\;\|\bm{x}\|_{1}.$ (5)
Additionally, we denote the set of at least $(\varepsilon,p)$-fair vectors as
$\mathcal{X}(\varepsilon,p)\;\triangleq\;\bigl\{\bm{x}\in\mathbb{R}_{\geqslant 0}^{n}:(1+\varepsilon\,D_{p})\,\|\bm{x}\|_{p}\leqslant\|\bm{x}\|_{1}\bigr\}.$ (6)
In the above definition, the factor $(1+\varepsilon D_{p})$ can be interpreted as the convex combination of the endpoints $1$ and $(D_{p}+1)$, with $\varepsilon$ as the non-negative multiplier. Furthermore, the constraint in (5) is convex for every $p\geqslant 2$. In particular:
  • For $p=2$, (5) reduces to a second‐order cone [10].
  • For $2<p<\infty$, it defines an $\ell_{p}$‐norm cone.
  • For $p=\infty$, it reduces to a collection of linear inequalities as follows:
    $\|\bm{x}\|_{1}\;\geqslant\;(1+\varepsilon D_{\infty})\,\|\bm{x}\|_{\infty}$
    $\;\Longleftrightarrow\;$ $\sum_{i=1}^{n}x_{i}\;\geqslant\;\left(1+n\varepsilon-\varepsilon\right)x_{i}\quad\forall\,i=1,\dots,n.$ (7)
For any fixed $\varepsilon$, the Definition 1 yields a countably infinite family of convex models (by varying the integer $p\geqslant 2$) for enforcing fairness in the distribution of $\bm{x}$. We now present a proposition that directly follows from the definition of $\mathcal{X}(\varepsilon,p)$ in (6).
Proposition 1.
For any integer $p\geqslant 2$ and and $\varepsilon_{1},\varepsilon_{2}\in[0,1]$ such that $\varepsilon_{1}>\varepsilon_{2}$, $\mathcal{X}(\varepsilon_{1},p)\subset\mathcal{X}(\varepsilon_{2},p)$. ∎
We now introduce a normalization argument to simplify notation, before stating and proving our main results.
Remark 1.
Normalization – Since all finite-dimensional norms are positively homogeneous of degree 1 [19], the inequality in (6), i.e.,
$\bigl(1+\varepsilon\,D_{p}\bigr)\,\|\bm{x}\|_{p}\;\leqslant\;\|\bm{x}\|_{1}$
is invariant under scaling of $\bm{x}$. Thus, it is convenient to normalize any non-zero vector [^1] The zero vector, $\bm{0}$, cannot be normalized and is trivially $(\varepsilon,p)$-fair for any $p\geqslant 2$ and $\varepsilon\in[0,1]$. Hence, throughout the rest of the article, we assume that $\bm{x}$ is always a non-zero vector. by its $\ell_{1}$-norm. To that end, define
$\Delta_{n}\;\triangleq\;\left\{\bm{x}\in\mathbb{R}_{\geqslant 0}^{n}:\sum_{i=1}^{n}x_{i}=1\right\},$ (8)
the standard probability simplex in $\mathbb{R}^{n}$. Normalizing any $\bm{x}\in\mathcal{X}(\varepsilon,p)$ with respect to its $\ell_{1}$-norm then yields
$\mathcal{Y}(\varepsilon,p)\;\triangleq\;\left\{\bm{x}\in\Delta_{n}:(1+\varepsilon\,D_{p})\,\|\bm{x}\|_{p}\leqslant 1\right\}.$ (9)
Observe that $\mathcal{Y}(\varepsilon,p)$ is exactly the image of $\mathcal{X}(\varepsilon,p)$ under the map $\bm{x}\mapsto\bm{x}/(\bm{e}^{\intercal}\bm{x})$; here $\bm{e}$ is a vector of 1s. Hence, any theoretical result for $\mathcal{X}(\varepsilon,p)$ can be transferred immediately to $\mathcal{Y}(\varepsilon,p)$ and vice-versa. All theoretical properties of $(\varepsilon,p)$-fairness in the subsequent section will be proved for the normalized set $\mathcal{Y}(\varepsilon,p)$.
III Theoretical Properties
We first establish the relationship between the $(\varepsilon,p)$-fairness of Definition 1 and the coefficient of variation of the vector $\bm{x}$. The coefficient of variation of $\bm{x}$ is a statistical measure of the dispersion of $\bm{x}$. This result provides the intuition for how varying $\varepsilon$ in (5) for any fixed integer $p\geqslant 2$ allows us to control of dispersion of $\bm{x}$.
III-A Bound of Coefficient of Variation
The first property we establish is that the $(\varepsilon,p)$‐fairness constraint (5) implies an upper bound on the coefficient of variation of $\bm{x}$, and that this bound tightens as $\varepsilon$ increases.
Proposition 2.
For any $\varepsilon\in[0,1]$, an integer $p\geqslant 2$, and $\bm{x}\in\mathcal{Y}(\varepsilon,p)$ defined in (9), the coefficient of variation of $\bm{x}$
$\mathrm{CV}(\bm{x})=\frac{\sqrt{\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\mu)^{2}}}{\mu},\quad\mu=\frac{1}{n},$
satisfies the following bound:
$\left(\mathrm{CV}(\bm{x})\right)^{2}\;\leqslant\;B_{p}(\varepsilon)\;\triangleq\;\frac{(D_{p}+1)^{2}}{(1+\varepsilon\,D_{p})^{2}}-1\,.$ (10)
Moreover, $B_{p}(\varepsilon)$ is strictly decreasing in $\varepsilon$ on $[0,1]$, with $B_{p}(0)=(D_{p}+1)^{2}-1$ and $B_{p}(1)=0$.
Proof.
For any $p\geqslant 2$, using norm equivalence $\mathcal{NE}(2,p)$ in (3) and (9), we have
$\|\bm{x}\|_{2}\;\leqslant\;\frac{D_{p}+1}{\sqrt{n}}\|\bm{x}\|_{p}\;\leqslant\;\frac{D_{p}+1}{\sqrt{n}}\left(\frac{1}{1+\varepsilon\,D_{p}}\right).$ (11)
On the other hand, since $n\mu=1$ and observing
$\frac{1}{n}\sum_{i=1}^{n}x_{i}^{2}=\frac{\|\bm{x}\|_{2}^{2}}{n},\quad\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\mu)^{2}=\frac{\|\bm{x}\|_{2}^{2}}{n}-\mu^{2},$
we obtain
$\left(\mathrm{CV}(\bm{x})\right)^{2}=\frac{\|\bm{x}\|_{2}^{2}-n\mu^{2}}{n\mu^{2}}=n\,\|\bm{x}\|_{2}^{2}-1.$
Substituting the bound on $\|\bm{x}\|_{2}$ from (11) yields
$n\,\|\bm{x}\|_{2}^{2}-1$ $\;\leqslant\;\frac{(D_{p}+1)^{2}}{(1+\varepsilon\,D_{p})^{2}}-1.$
Hence,
$\left(\mathrm{CV}(\bm{x})\right)^{2}\;\leqslant\;\frac{(D_{p}+1)^{2}}{(1+\varepsilon\,D_{p})^{2}}-1=B_{p}(\varepsilon).$
Finally, since $D_{p}>0$, one computes
$B_{p}^{\prime}(\varepsilon)\;=\;-\,\frac{2D_{p}(D_{p}+1)^{2}}{(1+\varepsilon\,D_{p})^{3}}\;<\;0.$
Hence, $B_{p}(\varepsilon)$ is strictly decreasing in $\varepsilon$ and $B_{p}(0)=(D_{p}+1)^{2}-1$, $B_{p}(1)=0$, as claimed. ∎
The strict monotonicity of $B_{p}(\varepsilon)$ in $\varepsilon$ implies that, as $\varepsilon$ increases, the upper bound on $\mathrm{CV}(\bm{x})$ decreases, thereby decreasing the dispersion of $\bm{x}$. This relationship was previously established for the case $p=2$ in [10].
In the forthcoming sections, we present the main theoretical results of this article on the equivalence or ordered relation between sets $\mathcal{Y}(p,\varepsilon)$ and correspondingly of $\mathcal{X}(p,\varepsilon)$ for varying values of $p$.
III-B Equivalence Condition
In this section, we give conditions under which requiring $\bm{x}$ to be at least $(\varepsilon,p)$‑fair is equivalent for every integer $p\geqslant 2$. We start with a key result that directly follows from the definition of $(\varepsilon,p)$-fairness.
Proposition 3.
Suppose that $\bm{e}_{i}$ is a unit vector with the $i$th coordinate set to $1$ and $\bm{e}$ is a vector of 1s, then for any integer $p\geqslant 2$, the following statements hold:
$\text{For any }i\in\{1,\dots,n\},~~\bm{e}_{i}\in\mathcal{Y}(\varepsilon,p)~\Longrightarrow~\varepsilon=0$ (12)
 $\text{ and }\varepsilon=1\iff\mathcal{Y}(\varepsilon,p)=\left\{\tfrac{1}{n}\bm{e}\right\}.$ (13)
Proof.
The proof follows from the observation in Section II that the term $(1+\varepsilon D_{p})$ is the convex combination of the two scalar values $1$ and $(D_{p}+1)$ with $\varepsilon$ as the factor. ∎
Theorem 1.
For any integer $p\geqslant 2$, and $\varepsilon\in\{0,1\}$, the sets of at least $(\varepsilon,p)$-fair vectors are equivalent, i.e.
$\mathcal{Y}(\varepsilon,p)=\mathcal{Y}(\varepsilon,2)\quad\forall~p>2.$
Proof.
We treat the two cases $\varepsilon=0$ and $\varepsilon=1$ separately. When $\varepsilon=0$, the inequality in (9) is trivially satisfied for any $\bm{x}\in\Delta_{n}$ and any $p\geqslant 2$. Hence,
$\mathcal{Y}(0,p)=\Delta_{n}=\mathcal{Y}(0,2)\quad\forall,p>2.$
When $\varepsilon=1$, Proposition 3 indicates that $\mathcal{X}(1,p)=\tfrac{1}{n}\bm{e}$ for every $p\geqslant 2$. ∎
Theorem 1 shows that at the two extreme fairness levels, the feasible set of vectors $\bm{x}$ is the same regardless of the choice of $p$.
III-C Strict Inclusion Property
Here, we prove a result that shows a strict inclusion property between $\mathcal{Y}(\varepsilon,p_{1})$ and $\mathcal{Y}(\varepsilon,p_{2})$ for any pair of integers $(p_{1},p_{2})$ such that $2\leqslant p_{1}<p_{2}$, and $\varepsilon$ in the open interval $(0,1)$. We begin by presenting a proposition that will be useful to prove the main theorem.
Proposition 4.
For any $\bm{x}\in\Delta_{n}\setminus\left\{\bm{e}_{1},\dots,\bm{e}_{n},\tfrac{1}{n}\bm{e}\right\}$,
$F(p)\triangleq\frac{1-\|\bm{x}\|_{p}}{D_{p}\|\bm{x}\|_{p}}\text{ is strictly decreasing in }p\geqslant 2.$ (14)
Proof.
See Appendix -A. ∎
Theorem 2.
For any pair of integers $(p_{1},p_{2})$ such that $2\leqslant p_{1}<p_{2}$ and $\varepsilon\in(0,1)$, $\mathcal{Y}(\varepsilon,p_{2})\subset\mathcal{Y}(\varepsilon,p_{1})$.
Proof.
For each $\bm{x}\in\Delta_{n}$ and integer $p\geqslant 2$, define
$\varepsilon_{p}(\bm{x})\;\triangleq\;\frac{1-\|\bm{x}\|_{p}}{D_{p}\,\|\bm{x}\|_{p}}.$ (15)
Since
$(1+\varepsilon\,D_{p})\,\|\bm{x}\|_{p}\leqslant 1,\iff\varepsilon\leqslant\varepsilon_{p}(\bm{x}),$
we have
$\bm{x}\in\mathcal{Y}(\varepsilon,p)\iff 0<\varepsilon\leqslant\varepsilon_{p}(\bm{x}).$
Therefore $\mathcal{Y}(\varepsilon,p_{2})\subset\mathcal{Y}(\varepsilon,p_{1})$ will follow once we show
$\varepsilon_{p_{1}}(\bm{x})>\varepsilon_{p_{2}}(\bm{x})~~\forall~~\bm{x}\in\Delta_{n}\setminus\left\{\bm{e}_{1},\dots,\bm{e}_{n},\tfrac{1}{n}\bm{e}\right\}.$ (16)
The excluded cases correspond to the boundary situation $\varepsilon\in\{0,1\}$, already treated in Proposition 3 and Theorem 1. For any such $\bm{x}\in\Delta_{n}\setminus\left\{\bm{e}_{1},\dots,\bm{e}_{n},\tfrac{1}{n}\bm{e}\right\}$, the proof is complete if we show that the map
$p\mapsto\varepsilon_{p}(\bm{x})=\frac{1-\|\bm{x}\|_{p}}{D_{p}\|\bm{x}\|_{p}}$
is strictly decreasing on $2,\infty]$ which is exactly the statement of Proposition 4. Figure 1 illustrates the idea graphically.
Figure 1: Graphical illustration for $\varepsilon_{p_{1}}(\bm{x})>\varepsilon_{p_{2}}(\bm{x})$.
Theorem 2 shows that the size of the feasible sets of requiring $\bm{x}$ to be at least $(\varepsilon,p)$-fair decreases with $p$ when $\varepsilon\in(0,1)$ and Theorem 1 shows that the feasible sets are equal for any value of integer $p\geqslant 2$ when $\varepsilon\in\{0,1\}$. Next, we consider an example application to illustrate the use of the models presented thus far in this article.
IV Multi-Agent Assignment Problem
In this section, we formulate a basic version of a multi-agent assignment problem (MAAP) and incorporate the fairness models introduced in Section II to control the dispersion of the total time required to complete the tasks across different agents. We start by introducing the relevant notation. Let $\mathcal{A}$ denote the set of agents and $\mathcal{J}$ the set of tasks. We assume that $|\mathcal{J}|>|\mathcal{A}|$. Each task $j\in\mathcal{J}$ must be completed by exactly one agent. For any agent $a\in\mathcal{A}$ and task $j\in\mathcal{J}$, let $t_{a,j}$ denote the time taken by agent $a$ to complete task $j$. The multi-agent assignment problem seeks to assign each task $j\in\mathcal{J}$ to one of the agents in $\mathcal{A}$ to minimize the total completion time. Furthermore, we introduce binary variables
$x_{a,j}=\begin{cases}1,&\text{if task }j\text{ is assigned to agent }a,\\ 0,&\text{otherwise.}\end{cases}$
Using these binary decision variables, a Mixed-Integer Linear Program (MILP) for the MAAP is as follows:
$\min:$~~ $\sum_{a\in\mathcal{A}}t_{a}$ (17a)
subject to: $t_{a}=\sum_{j\in\mathcal{J}}t_{a,j}\cdot x_{a,j}~~\forall a\in\mathcal{A},$ (17b)
$\sum_{a\in\mathcal{A}}x_{a,j}=1~~\forall j\in\mathcal{J},$ (17c)
$x_{a,j}\in\{0,1\}~~\forall a\in\mathcal{A},\;j\in\mathcal{J}.$ (17d)
Although the formulation in (17) is presented as an MILP, its constraint matrix is the node–arc incidence matrix of a bipartite graph, which is totally unimodular. Consequently, the linear programming relaxation has integral extreme points, and solving it via any linear programming solver yields the optimal integer solution [24]. Equivalently, one can apply the Hungarian algorithm directly to obtain the optimum in $O(|\mathcal{J}|^{3})$ time [25]. In contrast, a natural generalization in which each task carries a weight and each agent has a capacity constraint becomes NP‑hard (the generalized assignment problem) [26]. In this work, we restrict attention to the uncapacitated variant in (17), focusing on illustrating how $(\varepsilon,p)$‑fairness models control the dispersion of the processing times $t_{a}$ across agents in the simplest assignment problem, rather than a general variant of the MAAP.
IV-A Fairness-Constrained Multi-Agent Assignment Problem
We let $\bm{t}$ denote the vector of $t_{a}$ values over all agents $a\in\mathcal{A}$. The fairness‑constrained MAAP seeks to control the dispersion of the vector $\bm{t}$. This is enforced by requiring $\bm{t}$ to be at least $(\varepsilon,p)$‑fair for any $\varepsilon\in(0,1]$ and $p\geqslant 2$, as in (5). When $\varepsilon=0$, constraint (5) becomes trivial and the problem reduces to (17). The fairness constraint for fixed $p\geqslant 2$ and $\varepsilon\in(0,1]$ is
$(1+\varepsilon D_{p})\,\|\bm{t}\|_{p}\;\leqslant\;\sum_{a\in\mathcal{A}}t_{a}.$ (18)
Although the problem in (17) is solvable in polynomial time, adding (18) makes the resulting problem NP‑complete. Indeed, for two agents with $\varepsilon=1$, it reduces to the classical partition problem [27] that is known to be strongly NP-complete. As noted in Section II, when $p<\infty$, (18) is a convex constraint—specifically, an $\ell_{p}$‑norm conic constraint (for $p=2$, an $\ell_{2}$‑norm cone is referred to as a second‑order cone). Hence, the fairness‑constrained MAAP is a mixed‑integer convex program for $p<\infty$. When $p=\infty$, applying the reduction in (7) to (18) yields a mixed‑integer linear program that can be solved directly using off‑the‑shelf commercial or open‑source mixed‑integer linear programming solvers. For mixed‑integer convex optimization problems ($2\leqslant p<\infty$), constraint (18) can be lazily outer approximated by linear inequalities [28], enabling MILP solvers to obtain the solution. In the next section, we briefly present the outer approximation procedure for a general $\ell_{p}$-norm cone. We remark that the outer approximation algorithm is well-understood in the literature, and the presentation in the next section is only intended to make the article self-contained for the readers.
IV-B Outer Approximation of $\ell_{p}$-Norm Cone
We start with the definition of an $\ell_{p}$-norm cone [29] of dimension $(n+1)$ as
$P=\left\{(z,\bm{y})\in\mathbb{R}^{n+1}:z\geqslant\|\bm{y}\|_{p}\right\}.$ (19)
Note that the fairness constraint in (18) for any $p\geqslant 2$ can be represented in the form (19) by introducing an auxiliary variable. It is known in the literature that when convex cones are outer approximated by linear inequalities, it is both theoretically stronger and computationally more efficient to decompose a high-dimensional cone into several lower-dimensional cones and approximate each of those individually, rather than approximating the high-dimensional cone directly [30]. Converting a high-dimensional conic constraint into many low-dimensional cones is called “cone disaggregation” [30]. To that end, we present the steps to disaggregate the $(n+1)$-dimensional $\ell_{p}$-norm cone in (19) into $n$ three-dimensional cones plus a single linear constraint:
$z^{p}\;\geqslant\;\sum_{i=1}^{n}y_{i}^{p}~~\equiv~~z\;\geqslant\;\sum_{i=1}^{n}\frac{y_{i}^{p}}{z^{p-1}},$
$\equiv~~\left\{\,z=\sum_{i=1}^{n}Y_{i}\;\text{and}\;Y_{i}\,z^{p-1}\geqslant y_{i}^{p}~~\forall\,i\right\}$
$\equiv~~\left\{\,z=\sum_{i=1}^{n}Y_{i}\;\text{and}\;Y_{i}^{\frac{1}{p}}\,z^{1-\frac{1}{p}}\geqslant y_{i}~~\forall\,i\right\}.$
In the last equivalence, each constraint $Y_{i}^{\tfrac{1}{p}}\,z^{1-\tfrac{1}{p}}\geqslant y_{i}$ is convex for $2\leqslant p<\infty$ and in fact defines a three-dimensional power cone [29]:
$\mathcal{P}_{3}^{\alpha,1-\alpha}$ $\triangleq\left\{(v_{1},v_{2},v_{3})\in\mathbb{R}_{\geqslant 0}^{3}:v_{1}^{\alpha}v_{2}^{1-\alpha}\geqslant v_{3}\right\}.$ (21)
Hence, the cone in (19) can be rewritten as
$\left\{\,z=\sum_{i=1}^{n}Y_{i}\;\text{and}\;(Y_{i},z,y_{i})\in\mathcal{P}_{3}^{1/p,1-1/p}~~\forall\,i\right\},$ (22)
which comprises $n$ three-dimensional power cones and one linear equality. A linear outer approximation of each power cone is obtained by the first-order Taylor expansion at a point $(Y_{i}^{0},z^{0},y_{i}^{0})$ on its surface (i.e., satisfying the conic constraint at equality). Specifically, the linear outer approximation of $(Y_{i},z,y_{i})\in\mathcal{P}_{3}^{1/p,1-1/p}$ at $(Y_{i}^{0},z^{0},y_{i}^{0})$ is
$\left(\frac{1}{p}\right)\,y_{i}^{0}\,z^{0}\,Y_{i}\;+\;\left(1-\frac{1}{p}\right)\,y_{i}^{0}\,Y_{i}^{0}\,z\;\geqslant\;Y_{i}^{0}\,z^{0}\,y_{i}.$ (23)
In the algorithm to solve the mixed-integer convex optimization problem, we relax the $n$ three-dimensional power cones and iteratively add these linear outer approximations whenever the solution to the relaxed problem violates one of them. The required point on the surface of the power cone is computed by projecting the infeasible solution onto the corresponding power cone.
V Computational Results
This section provides empirical validation of the theoretical results in Section III using the formulations from Section IV. All experiments were run on 50 randomly generated instances of the MAAP, each with 5 agents and 250 tasks; the processing times $t_{a,j}$ were drawn uniformly at random. Both variants–with and without fairness constraints–were implemented in Kotlin and solved using CPLEX 22.1. The outer‑approximation algorithm of Section IV-B was realized via CPLEX’s lazy‑constraint callback. Source code for instance generation and both solution methods is available at https://github.com/abhay-singh11/FairTaskAllocation.
For each instance we solved the base assignment problem (17) augmented by the $(\varepsilon,p)$–fairness constraint (18) for
$p\in\{2,3,5,10,\infty\},\quad\varepsilon\in\{0,0.1,\dots,0.9,0.99,1.0\}.$
We found that all 50 instances were infeasible at $\varepsilon=1.0$, indicating that that perfect fairness cannot be enforced for the instances; we attribute this behaviour to the fact that $\{t_{a,j}\}$ are generated randomly. For this reason, we include $\varepsilon=0.99$ in our study: at this level, all instances remain feasible for every chosen $p$.
We let $\bm{t}^{*}(\varepsilon,p)$ denote the the optimal solution vector for the MAAP under fixed $(\varepsilon,p)$, and $\bm{t}^{*}(0)$ for the solution of the unconstrained problem (17). This notation is chosen to explicitly indicate that when $\varepsilon=0$ in (18), it is equivalent to imposing no fairness constraints on the MAAP.
V-A Dispersion Control
Here, we present the results for one representative instance out of the 50 randomly generated ones, since the other 49 exhibit the same qualitative behaviour. Figure 2 reports the coefficient of variation $\mathrm{CV}(\bm{t}^{*}(\varepsilon,p))$, computed as in Proposition 2. First, we observe that $\mathrm{CV}(\bm{t}^{*}(\varepsilon,p))$–the measure of dispersion–decreases monotonically with $\varepsilon$. Proposition 2 establishes that the upper bound on the coefficient of variation decreases in $\varepsilon$; Figure 2 confirms that the actual $\mathrm{CV}(\bm{t}^{*}(\varepsilon,p))$ likewise decreases. Next, we note the trend of diminishing returns: beyond $\varepsilon\approx 0.70$, the marginal reduction in dispersion is negligible, indicating that least fairness gains occur at moderate values of $\varepsilon$.
Figure 2: Coefficient of variation $\mathrm{CV}(\bm{t}^{*}(\varepsilon,p))$ versus $\varepsilon$ for different $p$ for one instance.
V-B Price of Fairness
The term “Price of Fairness” (PoF) is defined as the percentage increase in the objective value of the fairness‑constrained MAAP of Section IV-A relative to that of the unconstrained problem (17). This definition is standard in the literature [31, 10]. Formally,
$\mathrm{PoF}(\varepsilon,p)\;\triangleq\;\left(\frac{\bm{t}^{*}(\varepsilon,p)^{\intercal}\bm{e}}{\bm{t}^{*}(0)^{\intercal}\bm{e}}-1\right)\,\times 100\%.$
Figure 3: $\mathrm{PoF}(\varepsilon,p)$ versus $\varepsilon$ for various pp.
Figure 3 illustrates this trend for one representative instance; the other 49 instances show the same qualitative behavior. The increasing ordering $\mathrm{PoF}(\varepsilon,p_{1})\geqslant\mathrm{PoF}(\varepsilon,p_{2})$ for $p_{1}>p_{2}$ reflects the inclusion property of Theorem 2.
V-C Computation Times
The box‑plots of computation times are shown in Figure 4. Two trends stand out: (i) The computation time increases with $\varepsilon$. As $\varepsilon$ grows, the feasible region shrinks, making it harder for the solver to find and prove feasibility. (ii) For $p=\infty$, solve times are generally lower than for finite $p$. In that case, the fairness constraint reduces to a set of linear inequalities, whereas for $p<\infty$ the conic constraints incur additional overhead. Furthermore, when $p<\infty$, computation times are observed to increase with $p$.
(a) $\varepsilon=0.5$
(b) $\varepsilon=0.7$
(c) $\varepsilon=0.9$
(d) $\varepsilon=0.99$
Figure 4: Box-plot of computation times in seconds for all the 50 instances for $\varepsilon\in\{0.5,0.7,0.9,0.99\}$.
VI Conclusion
In this article, we present a countably infinite family of convex models–parameterized by $\varepsilon\in[0,1]$–to control the dispersion of a subset of decision variables in an optimization problem. The parameter $\varepsilon$ bounds dispersion: when $\varepsilon=0$, dispersion is unconstrained; when $\varepsilon=1$, dispersion is forced to zero. This family is indexed by integers $p\geqslant 2$, and all the models in the family are derived from first principles using the fundamental result of equivalence of finite-dimensional norms. For $p=2$ the model is a second‑order cone; for $2<p<\infty$ it is an $\ell_{p}$‑norm cone; and for $p=\infty$ it reduces to a set of linear inequalities.
We prove several theoretical properties that align with intuitive notions of fair solutions. We then evaluate the trade‑offs of these models on the MAAP. In particular, the $p=\infty$ variant is appealing both computationally–since the model is linear–and in terms of dispersion control, as increasing $\varepsilon$ yields a more gradual reduction in dispersion. Finally, these models can be integrated into optimization models in domains such as robotics, logistics, and transportation to control dispersion or enforce fairness of a vector of decision variables with minimal modifications to existing solution algorithms.
Acknowledgement
The authors thank Dirk Lauinger of the MIT Energy Initiative for insightful discussions on the potential for expanding on a family of fairness models.